L^{A}TEX

Anton Setzer^{∗}

7 October 2011

L^{A}TEX is a document preparation system for the TEX typesetting program. It offers programmable desktop publishing features and extensive facilities for automating most aspects of typesetting and desktop publishing, including numbering and cross-referencing, tables and figures, page layout, bibliographies, and much more. L^{A}TEX was originally written in 1984 by Leslie Lamport and has become the dominant method for using TEX; few people write in plain TEX anymore. The current version is L^{A}TEX 2ε. L^{A}TEX plays an important rôle in publishing scientific articles in Science.

E = mc^{2}  (1)
m = \frac{m_{0}}{\sqrt{1 − \frac{v^{2}}{c^{2}}}}  (2)

∗Dept. of Computer Science, Swansea University, Singleton Park, Swansea SA1 4PZ, UK. Email: a.g.setzer@swan.ac.uk. This article is based on the example in Wikipedia, http://en.wikipedia.org/wiki/LaTeX